OpenStudy (anonymous):

Center (7, 5), radius = 4

OpenStudy (aravindg):

The formula for a circle with center (h,k) and radius 4 is (x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

hi :) I suppose you're asking for the equation of a circle (x - j)^2 + (y - k)^2 = radius^2 there it is!

OpenStudy (anonymous):

Oh, aravindg said it first!

OpenStudy (anonymous):

but it should have been radius 'r' :)